18.01, Calculus, Problem Set 4

Problems to be collected, March 11: • §4.5: 2,3,6,11

• §4.6: 2,6,13

• §12.2: 2,4,9,11,22

• §12.3: 2,10,20,27

• Find the linearization of ^{3} \sqrt{1 + x} at x = 0.

• Use a linear approximation to estimate a) \sqrt{9.01}, b) ln 1.01.